It takes 8 minutes to load and 6 minutes to unload a machine. Inspection, packing, and travel between machines totals 6 minutes. Machines run automatically for 40 minutes. Operators cost $30 per hour; machines cost $55 per hour.

a. What is the maximum number of machines that can be assigned to an operator without creating machine idle time?

b. What is the number of machines that minimizes the cost?

c. If two machines are assigned to an operator, what will be the cost per unit produced?

d. If the cost per operator is unknown, for what range of values Co will the optimum assignment remain the same?
Answer #1

Time to load = 8 minutes

Time to unload = 6 minutes

Inspection, packing and travel time = 6 minutes

Run time = 40 minutes

Operator cost = $ 30 per hour

Machine cost = $ 55 per hour

1)

Total worker time per machine = 8+6+6 = 20 minutes

Total machine time per cycle = loading and unloading time + run time = 8+6+40 = 54 minutes

Number of machines that can be assigned to a worker without causing machine idle time = total machine time per cycle / total worker time per machine

= 54/20

= 2.7

2.7 is a fraction, if we round it up to 3, then total worker time per cycle (i.e. for 3 machines) will be 20*3 = 60 minutes, which is more than the cycle time for machine (54 minutes). This will cause machine idle time.

Therefore, maximum number of machines that can be assigned to an operator without creating machine idle time

= 2 machines

b)

If two machines are assigned to a operator, then the machine cycle time = 54 minutes

which means 2 units are produced every 54 minutes

Productivity per hour = 2 machines * 60/54

Total cost per hour = operator cost + machine cost = 30+2*55

= $ 140

Cost per unit produced = Total cost per hour / Productivity per hour

= 140 / (2*60/54)

= $ 63 per unit produced

c)

Optimum assignment is 2 machines per operator

The range of values of Co are determined by equating the cost of 2 machines with operator cost

So, maximum of Co is 2*55 = $ 110 per hour

Range of Co is from $ 0 to $ 110 per hour